# Modulo Sampling: Performance Guarantees in The Presence of Quantization

### v1

@@ VI Summary
This study analyzed the performance of a modulo ADC in the presence of a dithered quantization framework. We first considered the modulo ADC with 1-bit folding information and presented a recovery algorithm to unfold the modulo ADC samples. By using a dithered quantization framework for analysis, we showed that $\mathrm{OF}>3$ and $b>3$ are sufficient conditions for the modulo ADC with 1-bit folding information to achieve better quantization noise suppression when compared to a conventional ADC with the same number of bits used for amplitude quantization. Moreover, the MSE of modulo ADC decays faster than that of conventional ADC as $\mathrm{OF}$ is increased. When the 1-bit folding information is not available, we showed that OMP can be incorporated in the recovery algorithm to obtain the same MSE performance. However, the sufficient condition on the number of quantization bits is increased. We also analyzed the case of simultaneous acquisition of weak and strong signals occupying different frequency bands and demonstrated the superior performance of modulo ADC over conventional ADC. Numerical results are provided to substantiate the performance guarantees derived in this work.
@@ Appendix A Proof of Lemma 1
We first provide an upper bound for $|\mathcal{S}|$. Using [15, Equation 7], we have
@@
where the second inequality comes from the trival upper bound of the floor function, i.e., $\lfloor x\rfloor\leq x$. Meanwhile, $K=N-2\lfloor\frac{\rho N}{2}\rfloor\geq(1-\rho)N$. Hence, for $\mathbf{V}_{\mathcal{S}}$ to be a full column rank matrix, it suffices to show that
@@ Appendix B Proof of Theorem 1
We first bound the $\ell_{\infty}$-norm of the estimate of first-order difference of the modulo residue:
@@
where ${\Delta}\boldsymbol{\epsilon}_{\mathcal{S}}$ contains the values of $\Delta\epsilon[n]$ at the indices contained in $\mathcal{S}$. The third line follows because $\mathbf{V}_{\mathcal{S}}$ is full column rank by Lemma 1. The fourth line holds because $\epsilon[n]$ induced by the triangle dither $d[n]$ has amplitude support $\left(-\frac{3\lambda}{2^{b}},\;+\frac{3\lambda}{2^{b}}\right)$. Hence, $|\Delta\epsilon[n]|\leq\frac{6\lambda}{2^{b}}$ for all $n$.
@@
which is an explicit assumption of the theorem. Under this setting, $z[n]-\hat{z}[n]=0$. Consequently, from (II-B), the MSE becomes
@@
Since the MSE grows quadratically with $\lambda^{\prime}$, we should set the value of $\lambda^{\prime}$ as low as possible. Noting that $\lambda^{\prime}$ should satisfy (18), we obtain the desired result by setting $\lambda^{\prime}=\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}$.

### v2

@@ VI Conclusion
In this study, we analyzed the performance of a dithered modulo ADC. We first considered the case where a 1-bit folding information signal is available and proposed a recovery algorithm for unfolding the modulo samples. Our analysis showed that the conditions $\mathrm{OF}>3$ and $b>3$ are sufficient for achieving better quantization noise suppression compared to a conventional ADC with the same amplitude quantization bit budget. Additionally, we demonstrated that the MSE of the modulo ADC decreases more rapidly with increasing $\mathrm{OF}$ than that of a conventional ADC. Recognizing that folding information is often unavailable in practical implementations, we also developed and analyzed an orthogonal matching pursuit (OMP)-based reconstruction algorithm that does not require an additional folding information signal. The proposed OMP-based recovery method achieves the same MSE performance as that of the original unfolding scheme but the lack of 1-bit folding information penalizes the sufficient condition for $b$. More precisely, the sufficient condition for the oversampling factor is still $\mathrm{OF}>3$ but the amplitude quantization bit budget becomes $b>3+\log_{2}(\zeta)$, where $\zeta$ is the argument of the $\log_{2}(\cdot)$ operator in (16). We also analyzed the case of simultaneous acquisition of weak and strong signals occupying different frequency bands and demonstrated the superior performance of modulo ADCs over conventional ADCs. Numerical results are provided to substantiate the performance guarantees derived in this work.
@@ Appendix A Proof of Lemma 1
We first provide an upper bound for $|\mathcal{S}|$. Since $n_{0}$ and $n_{1}$ are chosen such that truncation error is negligible, we can use the results of [20] regarding the sparsity of the first-order difference of the modulo residue. Using [20, Equation 7], we have
@@
where the second inequality comes from the trivial upper bound of the floor function, i.e., $\lfloor x\rfloor\leq x$. Meanwhile, $K=N-2\lfloor\frac{\rho N}{2}\rfloor\geq(1-\rho)N$. Hence, for $\mathbf{V}_{\mathcal{S}}$ to be a full column rank matrix, it suffices to show that
@@ Appendix B Proof of Theorem 1
We first bound the $\ell_{\infty}$-norm of the estimate of the first-order difference of the modulo residue:
@@
where ${\Delta}\boldsymbol{\epsilon}_{\mathcal{S}}$ contains the values of $\Delta\epsilon[n]$ in the indices contained in $\mathcal{S}$. The third line follows because $\mathbf{V}_{\mathcal{S}}$ is the full column rank according to Lemma 1. The fourth line holds because $\epsilon[n]$ induced by the triangle dither $d[n]$ has amplitude support $\left(-\frac{3\lambda}{2^{b}},\;+\frac{3\lambda}{2^{b}}\right)$. Hence, $|\Delta\epsilon[n]|\leq\frac{6\lambda}{2^{b}}$ for all $n$.
@@
which is an explicit assumption of the theorem. In this setting, $z[n]-\hat{z}[n]=0$. Consequently, from (II-B), the MSE becomes
@@
Since MSE grows quadratically with $\lambda^{\prime}$, we should set the value of $\lambda^{\prime}$ as low as possible. Noting that $\lambda^{\prime}$ should satisfy (23), we obtain the desired result by setting $\lambda^{\prime}=\frac{c\cdot\|f(t)\|_{\infty}}{\mathrm{OF}-2}$ for some positive constant $1\leq c\leq\mathrm{OF}-2$.